# Exploring Pessimism and Optimism Dynamics in Deep Reinforcement Learning

### v1

@@ Abstract
Off-policy actor-critic algorithms have shown promise in deep reinforcement learning for continuous control tasks. Their success largely stems from leveraging pessimistic state-action value function updates, which effectively address function approximation errors and improve performance. However, such pessimism can lead to under-exploration, constraining the agent’s ability to explore/refine its policies. Conversely, optimism can counteract under-exploration, but it also carries the risk of excessive risk-taking and poor convergence if not properly balanced. Based on these insights, we introduce Utility Soft Actor-Critic (USAC), a novel framework within the actor-critic paradigm that enables independent control over the degree of pessimism/optimism for both the actor and the critic via interpretable parameters. USAC adapts its exploration strategy based on the uncertainty of critics through a utility function that allows us to balance between pessimism and optimism separately. By going beyond binary choices of optimism and pessimism, USAC represents a significant step towards achieving balance within off-policy actor-critic algorithms. Our experiments across various continuous control problems show that the degree of pessimism or optimism depends on the nature of the task. Furthermore, we demonstrate that USAC can outperform state-of-the-art algorithms for appropriately configured pessimism/optimism parameters.
@@ 1 Introduction
Deep reinforcement learning (RL) faces significant challenges when navigating the complex landscape of high-dimensional state spaces and non-linear state-action functions. The deadly triad of function approximation, off-policy learning, and bootstrapping compounds these challenges, leading to instability and poor sample efficiency [19, 27, 33]. One particularly critical consequence is the overestimation bias [29], where the estimated state-action values ($Q$-values) exceed their true counterparts [31, 32]. The overestimation bias is typically addressed through a pessimistic approach that use the minimum between two critics’ estimations [10, 12]. While effective in reducing overestimation, this approach can lead to pessimistic under-exploration [7], which limits the agent’s ability to explore and discover new policies. A natural remedy for pessimistic under-exploration could involve integrating optimism into actors’ exploration; however, the balance between pessimism and optimism is not straightforward, as this balance varies across tasks and evolves during the training process [20].
The core focus of this paper is to explore the dynamics between pessimism and optimism in deep RL, especially within actor critics. To address this dynamic, we introduce the novel framework termed Utility Soft Actor-Critic (USAC). Unlike binary approaches, USAC considers pessimism and optimism as points along a spectrum. Our aim is to address the overestimation bias while avoiding pessimistic under-exploration. In the next section, we explore the current literature on this topic and elaborate on our main contributions.
2 Exploring pessimism and optimism in deep reinforcement learning
The primary focus of this work is to understand the interplay between pessimism and optimism and their impact on overestimation bias and pessimistic under-exploration. In this section, we delve into related work addressing these challenges, exploring various strategies to overcome them. Moreover, we conclude this section by outlining our main contributions.
Addressing overestimation bias with pessimism
A pessimistic approach has proven effective in reducing the overestimation bias [^1] The overestimation bias stems from inherent noise and approximation errors introduced by the function approximations and is further amplified by bootstrapping [29, 31, 32]. – as demonstrated in Twin Delayed DDPG [^2] The Deep Deterministic Policy Gradient (DDPG) algorithm was proposed by Lillicrap et al. [17]. (TD3) [10] – it also introduces challenges related to pessimistic under-exploration [7]. Soft Actor-Critic (SAC) [12] addresses this challenge by integrating an entropy term into the actor’s policy decision-making process, thus making the actor stochastic. By incorporating the entropy, SAC encourages exploration and diversifies the actor’s actions, aiming to alleviate the adverse effects of pessimistic under-exploration. However, despite its effectiveness, entropy regularization alone do not resolve the underlying issues.
Exploring optimism to counter pessimistic under-exploration
An intuitive approach to tackle pessimistic under-exploration is based on a general form of the optimism in the face of uncertainty principle. Ciosek et al. [7] introduced the Optimistic Actor-Critic (OAC) algorithm, which adopts an optimistic approximate upper confidence bound on the $Q$-value function to enhance policy exploration. However, it’s crucial to note that aggressive exploration, as encouraged by optimistic algorithms like OAC, can be risky, particularly in environments where the size of overestimation is not well-understood [20].
Figure 1: Impact of pessimism/optimism levels on critic and actor. Top: Final returns. Bottom: Estimation error (i.e., true - estimation); positive values means underestimation, whereas negative values means overestimation.
Balancing between pessimism and optimism
@@ 2 Exploring pessimism and optimism in deep reinforcement learning
The balance of pessimism and optimism is nuanced and it varies across different tasks and develops throughout the training process [20]. The Tactical Optimism and Pessimism (TOP) algorithm, introduced by Moskovitz et al. [20], addresses this variability by dynamically balancing pessimism and optimism using multi-armed bandits [6, 16]. However, unlike algorithms such as OAC, which lean towards optimism, TOP learns to adjust the degree of pessimism or optimism based on the prevailing uncertainty. However, it uses the same level of pessimism/optimism for both critic and actor training, which limits its versatility.
USAC and our main contributions
Inspired by these insights, we see pessimism and optimism as points along a spectrum rather than binary opposites. While TOP is rooted in the TD3 algorithm, our approach stems from SAC. We introduce a novel framework called USAC, which quantifies the inherent uncertainty in learning through a utility function. Unlike binary approaches, USAC adapts its exploration strategy based on the critics’ distribution through uncertainty estimates of their $Q$-values. Moreover, USAC permits the fine-tuning of pessimism or optimism levels during critic or actor training. Through this methodology, USAC strives for a balanced pessimism-optimism trade-off that overcomes both overestimation bias and pessimistic under-exploration. Figure 1 illustrates the impact of varying levels of pessimism and optimism on both the critic and actor components. This indicates that optimal levels of pessimism and optimism can effectively address both overestimation (or underestimation) and pessimistic under-exploration, thereby enhancing the overall return. Through empirical evaluations and comparisons with state-of-the-art algorithms, we demonstrate the versatility and robustness of USAC across diverse continuous control tasks, paving the way for more adaptive and efficient learning algorithms in the field of deep RL.
3 Preliminary
@@
Markov decision processes (MDPs)
@@
Policies
Let $\Pi=\{\pi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})\}$ be the set of policies. The interaction between the agent and the MDP $\mathcal{M}$ under some policy $\pi\in\Pi$ progresses iteratively: At each time step $t\in\mathbb{N}$, the agent observes state $s_{t}\in\mathcal{S}$, chooses action $a_{t}\in\mathcal{A}$ based on the policy $a_{t}\sim\pi(\cdot|s_{t})$, receives a (bounded) reward $r_{t}:=r(s_{t},a_{t})$, and transitions to the next state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. Additionally, we define $p^{\pi}(s^{\prime},a^{\prime}|s,a)=p(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})$ as the one-step transition probability from $(s,a)$ to $(s^{\prime},a^{\prime})$.
Maximum entropy RL
The standard goal of RL is to find a policy $\pi$ that maximizes the expected sum of discounted rewards $J(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ with initial state $s_{0}\sim p_{0}$ [27, 3, 28]. To improve exploration, we will consider the more general maximum entropy RL framework [36, 11, 12, 25]. In this framework, one aims not only to maximize the expected (discounted) return but also to maximize the (discounted) entropy of the actions suggested by the stochastic policies. Formally, this is achieved by incorporating an entropy term, tempered by a parameter $\alpha>0$, which governs the relative importance of the return compared to the policy entropy. The maximum entropy objective function is defined as $J_{\alpha}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{t}+\alpha\mathcal{H}(\pi(\cdot|s_{t})))\right]$, where the policy entropy term $\mathcal{H}(\pi(\cdot|s))=-\mathbb{E}_{a\sim\pi(\cdot|s)}[\log\pi(a|s)]$.
@@
$T^{\pi}Q(s,a)=r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}[Q(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})]\,,$ (1)
i.e., $T^{\pi}Q^{\pi}(s,a)=Q^{\pi}(s,a)$ for any $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Actor-critic algorithms
One approach to learning the optimal policy is through the framework of actor-critic algorithms. In these algorithms, an agent simultaneously learns a policy (the actor), responsible for selecting actions that maximize expected return, and an estimation of the state-action value function (the critic), responsible for evaluating the policy’s quality through iterative updates. Formally, the critic $Q$ estimates the state-action value function $Q^{\pi}$ (derived by following policy $\pi$) by
@@
where $\tilde{Q}$ is (again) an estimate of the state-action value function $Q^{\pi}$, which can differ from both the critic $Q$ and the target critic $\bar{Q}$. In scenarios where the agent do a full Bellman backup without approximation errors, the critic $Q$, the target critic $\bar{Q}$, and the critic used to update the actor $\tilde{Q}$ are all identical to $Q^{\pi}$ [27, 12]. However, variations in $Q$, $\bar{Q}$, and $\tilde{Q}$ are often introduced to address the challenges we outlined in Section 2; we will explore some of these variations in more details below.
Exploring pessimism and optimism in actor-critic algorithms
A common strategy to address overestimation involves estimating the state-action value function using twin critics, along with associated target critics that are delayed versions of these critics [17, 18, 19, 10]. Specifically, let $Q_{1},Q_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ be the two critics and $\bar{Q}_{1},\bar{Q}_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ the two target critics that are delayed versions of the critics. The pessimistic strategy, employed by TD3 [10], SAC [12] and OAC [7], is to use the minimum of the two target critics $\{\bar{Q}_{k}\}_{k=1,2}$ to compute the target value of the critics in (3):
@@
This target is used to train both critics $Q_{1}$ and $Q_{2}$. Similarly, the same approach is employed to update the actor (in TD3 and SAC). Here, the critic to update the actor, $\tilde{Q}$ in (4), is determined by the minimum of the two critics $\{Q_{k}\}_{k=1,2}$:
@@
Ciosek et al. [7] noted that the pessimistic approach, described in 5 and 6, can lead to pessimistic under-exploration. To prevent this, they proposed embracing the optimism in the face of uncertainty principle in their OAC algorithm. However, this approach can potentially make exploration too aggressive, thereby negatively impacting overall performance [20]. Instead, the balance between pessimism and optimism should be tailored to the problem’s nature. Moskovitz et al. [20] suggested using the following target critic to achieve this balance:
@@
where the mean $\bar{\mu}$ and (unbiased) standard deviation $\bar{\sigma}$ are constructed using the two target critics $\bar{Q}_{1},\bar{Q}_{2}$; $\bar{\mu}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))$ and $\bar{\sigma}(s,a)=\sqrt{\sum_{k=1,2}(\bar{Q}_{k}(s,a)-\bar{\mu}(s,a))^{2}}$. The critic to update the actor, $\tilde{Q}$, can be defined in a similar manner from the two critics $\{Q_{k}\}_{k=1,2}$. Moskovitz et al. [20] classify 7 as optimistic when $\beta\geq 0$ and pessimistic when $\beta<0$. In the special case when $\beta=-1/\sqrt{2}$, 7 simplifies to taking the minimum of the two critics – as in 5 and 6. The TOP algorithm tries to adjust the balance between pessimism and optimism by using a multi-armed bandit algorithm to select $\beta$ from the set $\{-1,0\}$; here $\beta=-1$ corresponds to a pessimistic estimate and $\beta=0$ to an optimistic one. Although $\beta$ can vary over time using a bandit algorithm, TOP uses the same $\beta$ for both $\bar{Q}$ and $\tilde{Q}$.
4 Utility Soft Actor-Critic (USAC): framework and algorithms
Unlike the aforementioned algorithms, our framework decouples the levels of optimism and pessimism for the actor and critic. This novel approach allows us to be pessimistic in learning the critic while being optimistic in actor training. By doing so, we address the overestimation bias observed during critic training and overcome the problem of pessimistic under-exploration faced by the actor. Consequently, our approach promotes both a stable critic and an explorative actor simultaneously.
In this section, we establish the theoretical groundwork for exploring the balance between optimism and pessimism. We introduce the Utility Soft Actor-Critic (USAC) framework, which seamlessly integrates the uncertainty of critics into the actor learning process. Our framework guarantees policy improvement, as detailed in Section 4.1, while offering flexibility in navigating the optimism-pessimism trade-off. Moving on to Sections 4.2 and 4.3, we present the USAC algorithm and demonstrate how specific choices within our framework can emulate the optimism and pessimism strategies of algorithms like TD3 [10], SAC [12], OAC [7], and TOP [20]; in Section 5, we will explore these varying levels of optimism and pessimism across five MuJoCo environments. The pseudo-code of USAC is provided in Algorithm 1.
Before opening the entire discussion, we first introduce how to quantify the uncertainty of the $Q$-functions.
Quantifying the uncertainty of the $Q$-value distribution

### v2

# Improving Actor-Critic Training with Steerable Action-Value Approximation Errors

@@ Abstract
Off-policy actor-critic algorithms have shown strong potential in deep reinforcement learning for continuous control tasks. Their success primarily comes from leveraging pessimistic state-action value function updates, which reduce function approximation errors and stabilize learning. However, excessive pessimism can limit exploration, preventing the agent from effectively refining its policies. Conversely, optimism can encourage exploration but may lead to high-risk behaviors and unstable learning if not carefully managed. To address this trade-off, we propose Utility Soft Actor-Critic (USAC), a novel framework that allows independent, interpretable control of pessimism and optimism for both the actor and the critic. USAC dynamically adapts its exploration strategy based on the uncertainty of critics using a utility function, enabling a task-specific balance between optimism and pessimism. This approach goes beyond binary choices of pessimism or optimism, making the method both theoretically meaningful and practically feasible. Experiments across a variety of continuous control tasks show that adjusting the degree of pessimism or optimism significantly impacts performance. When configured appropriately, USAC consistently outperforms state-of-the-art algorithms, demonstrating its practical utility and feasibility.
[…]
6112
@@ 1 Introduction
Deep reinforcement learning (RL) faces significant challenges when navigating the complex landscape of high-dimensional state spaces and non-linear state-action functions, especially in continuous control tasks. The deadly triad of function approximation, off-policy learning, and bootstrapping compounds these challenges, often leading to instability and poor sample efficiency [21, 30, 36]. One particularly critical consequence is overestimation bias [32], where the estimated action-values ($Q$-values) exceed their true counterparts [34, 35].
Overestimation bias in off-policy RL can arise from multiple sources. A well-studied source is the maximization operator in temporal-difference learning, which systematically drives estimates upward by preferring positive noises. It is typically addressed by min-clipping [12, 14], where the minimum of two critic networks is used, though this introduces a systematic underestimation bias [34]. Other contributors include sampling error in batch temporal-difference learning [24] and reuse bias, which arises when the same data are repeatedly used for updates [39].
Although the remedies reduce overestimation, they often swing too far toward pessimism, producing overly pessimistic policy evaluations discouraging exploration. This phenomenon, known as pessimistic under-exploration [9], limits effective learning in actor-critic methods, necessitating methods that can balance optimism and pessimism more systematically. However, the interaction between critic errors and resulting actor behavior follows complex dynamics in online deep RL [22], and remains poorly understood theoretically.
We posit that the impact of the interplay between optimism and pessimism on the learning profiles of actor-critic pipelines can be explained by a simple and intuitive pair of hyperparameters. We introduce a new framework that puts the action-value predictions of a pair of critics into a utility-theoretic perspective, which considers optimism and pessimism as points along a spectrum. This framework is practical as the hyperparameters controlling optimism and pessimism are interpretable and can be tuned for different tasks. The commonplace min-clipping boils down to only one of infinitely many possibilities within this spectrum. The descriptive value of this spectrum can be of the interest to theoreticians. It can also be used by practitioners as a hyperparameter to improve performance via grid-search.
2 Prior art of orchestrating optimism and pessimism
The state-of-the-art suggests the following three strategies to counteract the under-exploratory effects of pessimistic policy evaluation:
(i) Maximum-entropy policy search. Algorithms such as Soft Actor-Critic (SAC) [14] maintain exploration by incentivizing an entropy increase in the policy distribution. This incentive is added to the reward function in the Bellman target calculation and is accounted for also during actor training. In effect, this is a random exploration scheme that perturbs the action selection independently from an estimation of the accuracy of the action-value approximations. This disconnect manifests itself as the brittleness of the algorithm on the entropy coefficient [15].
(ii) Optimistic exploration. Many theoretically-grounded approaches for model-driven exploration adopt the Optimism in the Face of Uncertainty (OFU) principle [1, 2]. Ciosek et al. [9] implemented this principle into a deep actor-critic pipeline with their seminal Optimistic Actor-Critic (OAC) algorithm. OAC adopts an optimistic approximate upper confidence bound on the $Q$-value function to enhance policy exploration while using a pessimistic lower-confidence bound for critic training. The algorithm is computationally expensive. Furthermore, the effectiveness of its exploration is limited by the approximation errors incurred during confidence set calculation.
@@ 2 Prior art of orchestrating optimism and pessimism
(iii) Dynamic exploration. The balance of optimism and pessimism is nuanced, and it varies across different tasks and develops throughout the training process [22]. The Tactical Optimism and Pessimism (TOP) algorithm, introduced by Moskovitz et al. [22], addresses this variability by dynamically balancing optimism and pessimism multi-armed bandits [8, 18]. The approach assumes a Gaussian distributed action-value function and it tunes the degree of exploration by playing a bandit game on a predetermined grid of quantiles. The TOP algorithm uses the same level of pessimism/optimism for both critic and actor training in order to maintain the feasibility of the bandit game, which limits the versatility of the algorithm. Furthermore, as the algorithm is designed upfront for on-the-fly tuning, it does not permit an interpretable analysis of pessimism and optimism in environments of interest.
We propose a simpler alternative: treating the critic pairs as samples from a parametric distribution over action-value functions. For example, when using a Laplace distribution as our main focus in the paper, embedding it within a utility function allows us to express the degree of critic pessimism and actor optimism on an interpretable scale between 1-1 and +1+1. We demonstrate the expressive and instrumental power of this framework within a SAC extension due to the wide adoption of the model family. We refer to the resulting algorithm as a Utility Soft Actor Critic (USAC). However, our framework is equally applicable to alternative families such as deterministic policy gradient variants [19, 12]. As shown in Figure˜1, our framework reveals how the specific demands of an environment for pessimistic policy evaluation and optimistic policy search influence each other. The left panel illustrates the impact of different combinations of pessimism and optimism on final returns, where a coarse grid-search can yield improvements exceeding 30% in several cases. The right panel shows the corresponding effect on value estimation bias (true value minus estimated value), indicating whether the policy tends to under- or over-estimate returns.
Figure 1: Impact of optimism and pessimism levels on actor–critic performance on Walker2d environment. Left: Final returns for different combinations of pessimism and optimism. Right: Value estimation bias (i.e., true – estimated value); positive values indicate underestimation, negative values indicate overestimation
3 Preliminaries
@@
Markov decision processes (MDPs). An MDP can be represented by the tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},p,p_{0},r,\gamma\rangle$ [26]. Here, $\mathcal{S}$ and $\mathcal{A}$ denote the continuous state and action spaces. The function $p(s^{\prime}|s,a)$ for $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ represents the transition probability from the current state $s$ and action $a$ to the next state $s^{\prime}$. We consider the model-free setup where this function is not known and will not be learned. The initial state distribution is denoted by $p_{0}\in\mathcal{P}(\mathcal{S})$, while $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,B_{r}]$ represents the bounded reward function with $B_{r}>0$, and $\gamma\in[0,1]$ stands for the discount factor.
Policies. Let $\Pi=\{\pi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})\}$ be the set of policies. The interaction between the agent and the MDP $\mathcal{M}$ under some policy $\pi\in\Pi$ progresses iteratively. At each time step $t\in\mathbb{N}$, the agent observes state $s_{t}\in\mathcal{S}$, chooses action $a_{t}\in\mathcal{A}$ based on the policy $a_{t}\sim\pi(\cdot|s_{t})$, receives a (bounded) reward $r_{t}:=r(s_{t},a_{t})$, and transitions to the next state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. Additionally, we define $p^{\pi}(s^{\prime},a^{\prime}|s,a)=p(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})$ as the one-step transition probability from $(s,a)$ to $(s^{\prime},a^{\prime})$.
Maximum entropy RL. The standard goal of RL is to find a policy $\pi$ that maximizes the expected sum of discounted rewards $J(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ with initial state $s_{0}\sim p_{0}$ [30, 5, 31]. To improve exploration, we will consider the more general maximum entropy RL framework [40, 13, 14, 28]. In this framework, one aims not only to maximize the expected (discounted) return but also to maximize the (discounted) entropy of the actions suggested by the stochastic policies. Formally, this is achieved by incorporating an entropy term, tempered by a parameter $\alpha>0$, which governs the relative importance of the return compared to the policy entropy. The maximum entropy objective function is defined as $J_{\alpha}(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}(r_{t}+\alpha\mathcal{H}(\pi(\cdot|s_{t})))\right]$}, where the policy entropy term is defined as $\mathcal{H}(\pi(\cdot|s))=-\mathbb{E}_{a\sim\pi(\cdot|s)}[\log\pi(a|s)]$.
@@
that is, $T^{\pi}Q^{\pi}(s,a)=Q^{\pi}(s,a)$ for any $(s,a)\in\mathcal{S}\times\mathcal{A}$.
Actor-critic algorithms. Performing continuous control with deep RL necessitates a policy-gradient approach applicable to online updates. The actor-critic design is an optimal and widely adopted framework for this purpose. In these algorithms, an agent simultaneously learns a policy (the actor), responsible for selecting actions that maximize expected return, and an estimation of the action-value function (the critic), responsible for evaluating the policy’s quality through iterative updates. Formally, the critic $Q$ is trained to predict the action-value function $Q^{\pi}$ (derived by following policy $\pi$) by temporal difference learning
@@
where $\tilde{Q}$ is (again) an estimate of the action-value function $Q^{\pi}$, which can differ from both the critic $Q$ and the target critic $\bar{Q}$. In scenarios where the agent does a full Bellman backup without approximation errors, the critic $Q$, the target critic $\bar{Q}$, and the critic used to update the actor $\tilde{Q}$ are all identical to $Q^{\pi}$ [30, 14]. However, variations in $Q$, $\bar{Q}$, and $\tilde{Q}$ are often introduced to address the challenges outlined in the introduction. We will explore some of these variations in more detail below.
The sources of optimism and pessimism in actor-critics. A common strategy to address overestimation involves estimating the action-value function using twin critics, along with associated target critics that are delayed versions of these critics [19, 20, 21, 12]. Specifically, let $Q_{1},Q_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ be the two critics and $\bar{Q}_{1},\bar{Q}_{2}:\mathcal{S}\times\mathcal{A}\rightarrow{\mathbb{R}}$ the two target critics that are delayed versions of the critics. The pessimistic strategy, employed by TD3 [12], SAC [14] and OAC [9], is to use the minimum of the two target critics $\{\bar{Q}_{k}\}_{k=1,2}$ to compute the target value of the critics in (3):
@@
This target is used to train both critics $Q_{1}$ and $Q_{2}$. Similarly, the same approach is employed to update the actor (in TD3 and SAC). Here, the critic used to update the actor, $\tilde{Q}$ in (4), is determined by the minimum of the two critics $\{Q_{k}\}_{k=1,2}$:
@@
Ciosek et al. [9] noted that the pessimistic approach, described in ˜5 and 6, can lead to pessimistic under-exploration. To prevent this, they proposed embracing the optimism in the face of uncertainty principle in their OAC algorithm. However, this approach can potentially make exploration too aggressive, thereby negatively impacting overall performance [22]. Instead, the balance between optimism and pessimism should be tailored to the problem’s nature. Moskovitz et al. [22] suggested using the following target critic to achieve this balance:
@@
where the mean $\bar{\mu}$ and (unbiased) standard deviation $\bar{\sigma}$ are constructed using the two target critics $\bar{Q}_{1},\bar{Q}_{2}$:
$\bar{\mu}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))$
and
$\bar{\sigma}(s,a)=\sqrt{\sum_{k=1,2}(\bar{Q}_{k}(s,a)-\bar{\mu}(s,a))^{2}}.$
The critic to update the actor, $\tilde{Q}$, can be defined in a similar manner from the two critics $\{Q_{k}\}_{k=1,2}$. Moskovitz et al. [22] classify ˜7 as optimistic when $\beta\geq 0$ and pessimistic when $\beta<0$. In the special case when $\beta=-1/\sqrt{2}$, ˜7 simplifies to taking the minimum of the two critics as in ˜5 and 6. The TOP algorithm tries to adjust the balance between optimism and pessimism by using a multi-armed bandit algorithm to select $\beta$ from the set $\{-1,0\}$; here $\beta=-1$ corresponds to a pessimistic estimate and $\beta=0$ to an optimistic one. Although $\beta$ can vary over time using a bandit algorithm, TOP uses the same $\beta$ for both $\bar{Q}$ and $\tilde{Q}$.
4 A utility-theoretic framework to make the action-value biases steerable
Our key contribution is a new theoretical framework to quantify the degree of optimism or pessimism assumed by an actor-critic algorithm that performs policy evaluation via twin critics using two interpretable scalars, one for the critic and one for the actor. We achieve this by assuming that the action-value function follows a parametric distribution and the twin critics are two samples from this distribution. The resulting utility function yields a form that enables describing the optimism and pessimism levels via scalar risk coefficients that take the commonplace min-clipping method as a special case and provides an effortless generalization of it by simple grid-search.